OpenStudy (anonymous):

Can someone please show me the steps on how to solve this? x - y = 16 2x + y = 2

OpenStudy (anonymous):

add the first equation to the second in a column

OpenStudy (anonymous):

\(x+2x=3x\) , \(-y+y=0\) and \(16+2=16\) so you get \[3x=18\] solve for \(x\)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw once you have \(x\) it is easy to find \(y\) right?

OpenStudy (anonymous):

I think so hold on please :)

OpenStudy (anonymous):

What do I do from there? I know x=6 but how do I go into the next step?

OpenStudy (anonymous):

Oh wait I got it! Thanks so much for the help!!